### v1

# Evaluating Vision–Language and Large Language Models for Automated Student Assessment in Indonesian Classrooms

Abstract
Although vision-language and large language models (VLM and LLM) offer promising opportunities for AI-driven educational assessment, their effectiveness in real-world classroom settings, particularly in underrepresented educational contexts, remains underexplored. In this study, we evaluated the performance of a state-of-the-art VLM and several LLMs on 646 handwritten exam responses from grade 4 students in six Indonesian schools, covering two subjects: Mathematics and English. These sheets contain more than 14K student answers that span multiple choice, short answer, and essay questions. Assessment tasks include grading these responses and generating personalized feedback. Our findings show that the VLM often struggles to accurately recognize student handwriting, leading to error propagation in downstream LLM grading. Nevertheless, LLM-generated feedback retains some utility, even when derived from imperfect input, although limitations in personalization and contextual relevance persist.
1 Introduction
Vision–language models (VLMs) Liu et al. (2023, 2024b); Steiner et al. (2024) and large language models (LLMs) Touvron et al. (2023a); Team (2024); Team et al. (2024); OpenAI et al. (2024) have demonstrated impressive reasoning capabilities Wang et al. (2023); Wei et al. (2022), including solving complex academic tasks such as university-level physics Yeadon and Hardy (2024) and competition-grade mathematics problems Zhang et al. (2024). These advancements have driven growing interest in applying such models to education. Common areas of application include automated grading Chiang et al. (2024), teaching support Hu et al. (2025), feedback generation Morris et al. (2023), and content creation Westerlund and Shcherbakov (2024).
Figure 1: AI-powered assessment using VLM and LLM.
However, most VLM and LLM-based educational tools have been developed with English-speaking contexts in mind Lee and Zhai (2025); Yancey et al. (2023), limiting their relevance and usability in non-English-speaking regions, particularly in rural areas in Indonesia. Ensuring socio-cultural relevance is essential: effective deployment requires adaptation to local curricula, languages, and cultural norms, rather than relying on a one-size-fits-all approach. Moreover, the shortage of qualified teachers in rural areas highlights the importance of prioritizing AI integration in underserved regions, rather than concentrating development efforts solely in high-resource, Global North contexts Jin et al. (2025); Kristiawan et al. (2024).
In this study, we address the contextual challenges of applying AI-powered assessment tools in non-English speaking and under-resourced settings by collecting real-world student assessment data from primary schools in the form of handwritten responses. This design is motivated by two practical considerations. First, many schools, especially in rural areas, lack consistent access to digital devices, highlighting the need for AI systems that function effectively in low-tech environments. Second, using handwritten responses helps reduce the risk of academic dishonesty, such as students who rely on AI tools to generate answers. Assessments were conducted in Indonesian for the mathematics subject, while the responses to the English subject were written in English, reflecting the language of instruction for each subject.
Our contributions are as follows: (1) We release a dataset of 646 handwritten student answer sheets (with over 14K answers) collected from six primary schools in Indonesia—three from rural areas and three from urban areas. The assessments cover Grade 4 Mathematics and English, with questions and scoring guidelines developed by experienced teachers. All student responses were manually transcribed and graded by professional teachers. [^1] To ensure ethical use and protect student privacy, all personally identifiable information (e.g., student names, grade levels, and school names) has been removed. (2) We introduce a multimodal pipeline that integrates vision–language models (VLMs) and large language models (LLMs), as illustrated in Figure 1. We compare several state-of-the-art models for grading student answers and find that GPT-4o with vision input achieves the highest accuracy and feedback quality. (3) We conduct a manual evaluation of LLM-generated feedback in Indonesian and find that, even when based on imperfect input (e.g., OCR errors), the feedback tends to be clear and factually correct. However, personalization and helpfulness remain notable areas of concern.
2 Related Work
Previous studies have investigated the use of LLMs as graders for student assignments and exams. For example, Chiang et al. (2024) used GPT-4 to automatically grade 1,028 student essays in a university-level course titled Introduction to Generative AI. Their findings suggest that LLM-based graders were generally well accepted by students; however, the models occasionally did not follow the grading rubric. In a related study, Yancey et al. (2023) used GPT-3.5 and GPT-4 to score essays in a high-stakes English proficiency test, demonstrating that LLM-generated scores can achieve high agreement with human raters.
Stahl et al. (2024) used Mistral (Jiang et al., 2023) and LLaMA-2 (Touvron et al., 2023b) to assess English student essays and generate feedback, finding that scoring accuracy had limited influence on student’s perceived usefulness of the feedback. Similarly, Morris et al. (2023) applied a Longformer-based language model (Botarleanu et al., 2022) to generate formative feedback on student-written summaries of English textbooks.
Unlike these prior studies, our work focuses on handwritten responses from grade 4 primary school students in Indonesia, covering both English and mathematics. We also evaluate a complete multimodal pipeline that integrates a VLM for handwriting recognition and LLMs for grading and feedback generation—introducing new challenges related to noisy input, multilingual content, and real-world constraints in low-tech, underrepresented classroom settings.
3 Dataset Construction
Assessment Design
We developed assessment instruments for grade 4 primary school students in two subjects: Mathematics and English. The items were designed from scratch based on a thorough analysis of the national curriculum and corresponding learning objectives. Each subject assessment consisted of 10 multiple-choice questions (MCQs), 10 short-answer questions, and 2 essay questions. All items were created by experienced senior subject teachers—an English teacher and a Math teacher—each with over 10 years of classroom experience and a Master’s degree in Education. In addition to writing the assessment items, these teachers developed detailed scoring rubrics for the short-answer and essay questions, as well as answer keys for the MCQs. Standardized answer sheets were also prepared to collect student responses.
Data Collection
Data collection was carried out in six primary schools, evenly divided between rural (Sumatra and Nusa Tenggara Islands) and urban (Java Island) settings. Each classroom included approximately 20 to 30 students. For both subjects, students followed a structured sequence consisting of a pre-test, lesson, and post-test. Students had up to 30 minutes to complete their answers on a standardized answer sheet.
In total, we collected 646 handwritten answer sheets from these assessments, comprising both pre-tests and post-tests. Of these, 414 were collected from urban schools and 232 from rural schools. The disparity in sample size between urban and rural areas is primarily due to larger class sizes typically found in urban schools compared to their rural counterparts.
4 Experiment
Overall Pipeline
Figure 1 illustrates our pipeline, which begins with a vision–language model (VLM) that performs optical character recognition (OCR) to extract handwritten student responses from scanned answer sheets. The extracted text is then structured into a JSON format and passed to a large language model (LLM), along with the answer key and a teacher-defined rubric. For multiple-choice questions, we apply string matching. For short-answer and essay questions, we run the LLM separately for each question, providing the student’s response, the corresponding answer key, and the assessment rubric. To generate personalized feedback, we provide the LLM with all of the student’s responses, the answer key, the assigned weights, and the rubric.
Model
For OCR, we use GPT-4o OpenAI et al. (2024), alongside a gold-standard transcription manually parsed by teachers. For automatic scoring, we compare the performance of GPT-4o, Llama-3.1-Instruct (70B) Touvron et al. (2023b), Qwen2.5-Instruct (72B) Team (2024), and Deepseek-Chat (671B) Liu et al. (2024a). For generating personalized feedback, we rely on the scoring results produced by GPT-4o and generate two versions of feedback using GPT-4o and Deepseek-Chat. All prompts and decoding hyperparameters used are provided in the Appendix.
Evaluation
Each answer sheet image was manually transcribed and scored by professional teachers. We compared the LLM-generated scores against these gold-standard scores across three question types: multiple-choice, short-answer, and essay, using mean absolute error (MAE) as the evaluation metric. For personalized feedback, we conducted a manual evaluation covering four aspects—Correctness, Personalization, Clarity, and Educational Value/Helpfulness—rated on a 1–5 scale, where 1 indicates the lowest quality. [^2] This evaluation was carried out by an experienced educator with a Master’s degree in teaching. The evaluation guidelines and definitions for each aspect are provided in the Appendix.
English Math
Model M S E Total M S E Total
OCR by GPT4o
GPT4o 2.8 14.6 5.6 11.7 2.3 16.3 1.5 8.2
Llama 3.1 (70B) 2.8 18.7 9.3 14.5 2.3 10.6 27.5 2.2
Qwen2.5 (72B) 2.8 14.9 16.6 14.7 2.3 19.1 5.8 7.1
Deepseek (671B) 2.8 12.6 9.8 11.9 2.3 22.8 6.7 8.1
OCR by Human
GPT4o 0.0 9.2 2.7 7.9 0.0 2.9 5.7 1.5
Llama 3.1 (70B) 0.0 14.4 2.3 11.6 0.0 9.8 19.1 10.3
Qwen2.5 (72B) 0.0 8.4 3.8 9.2 0.0 5.5 8.7 3.3
Deepseek (671B) 0.0 4.4 1.5 6.8 0.0 5.9 8.5 0.8
Table 1: Mean absolute error (MAE) for scoring English and Math, calculated separately for multiple-choice (M), short-answer (S), essay (E), and the total score. Lower values indicate better performance; bolded numbers represent the best results. Scores for each component range from 0 to 100.
Model Correctness Personalization Clarity Helpfulness
English
GPT-4o 4.00 3.96 3.64 3.60
Deepseek 3.96 3.88 4.04 3.96
Math
GPT-4o 3.84 3.72 3.92 3.68
Deepseek 3.88 2.96 4.00 2.92
Table 2: Human evaluation by expert teachers on personalized feedback, using a rating scale from 1 to 5, where 1 indicates the lowest score.
5 Result and Analysis
Main Result
Table 1 presents the performance of the LLMs selected in three types of questions: multiple choice, short answer, and essay. When using GPT-4o to extract student responses via OCR, we observe that most model-generated scores are generally competitive. Among them, GPT-4o produces scores that align most closely with human grading for essay questions, achieving the lowest MAE in both English (5.6) and Math (1.5). In contrast, LLaMA-3.1–70B and Qwen-2.5–72B are less reliable, with scores deviating more significantly from human judgments. Short-answer questions remain the most challenging to evaluate: even the best performing model in this category, LLaMA-3.1-7B for Math, still shows a relatively high MAE of 10.6, indicating a notable gap from human-level accuracy.
However, the results differ when human effort is involved in the OCR task. Most scores become better overall, with Deepseek-chat and GPT-4o emerging as the top-performing models. Deepseek-chat shows strong performance in English (MAE of 4.4 for short answers and 1.5 for essays), while GPT-4o performs best in Math, with only a 2.9 difference in short answers and 5.7 in essays. It is worth noting that MCQ scores remain at 0, as basic string matching is sufficient due to the exact nature of the answers. The impact of OCR performance on LLM scoring is further discussed in Section 5.
Human Evaluation on Personalised Feedback
Table 2 presents the results of a human evaluation on personalized feedback quality, rated by expert teachers across four dimensions: Correctness, Personalization, Clarity, and Helpfulness (scale 1–5, with scores below 3 considered poor). For English, GPT-4o slightly outperforms Deepseek in correctness and personalization, while Deepseek leads in clarity and helpfulness. In Math, Deepseek shows strong clarity and correctness but performs poorly in personalization and helpfulness, with both scores falling below 3. GPT-4o, on the other hand, maintains more balanced performance across all dimensions.
English Math
Model M S E Total M S E Total
Urban
GPT4o 0.0 2.4 7.2 0.8 0.0 5.8 7.6 2.4
Llama 3.1 (70B) 0.0 7.7 2.9 2.7 0.0 10.3 30.0 10.4
Qwen2.5 (72B) 0.0 1.9 1.3 0.5 0.0 7.6 10.7 3.9
Deepseek (671B) 0.0 1.3 3.5 1.5 0.0 5.6 9.9 1.0
Rural
GPT4o 0.0 21.2 5.2 23.1 0.0 2.5 2.2 0.3
Llama 3.1 (70B) 0.0 26.1 11.4 26.9 0.0 8.8 23.1 9.7
Qwen2.5 (72B) 0.0 19.8 12.5 24.3 0.0 1.7 5.0 2.1
Deepseek (671B) 0.0 14.2 10.1 21.2 0.0 6.4 5.9 0.6
Table 3: Analysis of mean absolute errors (MAE) for scoring English and Math across urban and rural settings, calculated separately for multiple-choice (M), short-answer (S), essay (E), and total scores. The OCR results used in this analysis were obtained through human transcription. Lower values indicate better performance; bolded values represent the best results. Each component is scored on a 0–100 scale.
Urban vs. Rural Performance Analysis
Given the significant educational disparities between rural and urban areas, we evaluated the performance of the model in these two settings. To isolate the analysis of LLM scoring capabilities, we use only the human-transcribed OCR results, eliminating recognition errors.
Table 3 presents the MAE scores for English and Math, separated by question type: multiple choice (M), short answer (S), essay (E), and total scores. The results indicate that English MAEs are generally higher in rural settings than in urban settings across all models. For example, GPT-4o achieves a total MAE of only 0.8 in urban English, but this rises sharply to 23.1 in the rural setting. This discrepancy suggests that LLMs may struggle more in interpreting free-form responses from rural students, possibly due to variations in writing style and grammar. In contrast, MAEs for Math tend to be slightly lower in rural areas, although the differences are less pronounced. This may be attributed to the nature of Math questions, which often involve numerical reasoning and have more deterministic answers, reducing ambiguity in scoring.
OCR Performance Analysis
Given the differences in MAE between the GPT-4o OCR outputs and human transcription shown in Table 1, we further analyze the OCR performance of GPT-4o and evaluate the extent to which recognition errors propagate to the subsequent scoring. For this analysis, we use exact string matching to assess accuracy on multiple choice and short answer questions, and compute ROUGE-L Lin (2004) scores to compare GPT-4o and human transcriptions for essay questions.
Table 4 shows that the OCR performance is generally higher for English than for Math. Within English, responses from urban students yield higher exact match and ROUGE-L scores compared to those from rural students, possibly due to differences in handwriting clarity or writing conventions. For Math, the OCR accuracy is overall lower than that of English, but the performance gap between urban and rural settings is less pronounced. This suggests that while English responses may be more affected by region-specific handwriting variability, Math responses, often more structured and numerical, are comparatively stable across regions.
English Math
Area EM(M) EM(S) RL(E) EM(M) EM(S) RL(E)
Urban 82.1 67.1 60.3 62.3 23.3 21.0
Rural 71.7 61.8 60.1 62.5 27.9 24.8
All 78.5 65.3 60.2 62.4 24.9 22.3
Table 4: OCR-based performance (GPT-4o) across Urban, Rural, and All settings for English and Math: EM = exact match, RL = ROUGE-L F1, M = multiple choice, S = Short Essay, E = Essay.
6 Conclusion
In this work, we present a real-world implementation of vision–language model (VLM) and large language models (LLMs) for student assessment in underrepresented regions—specifically, rural and urban areas of Indonesia—focusing on primary school subjects in Math and English. Our results show that GPT-4o and Deepseek (671B) perform competitively in matching teacher-assigned scores across multiple-choice, short-answer, and essay formats. For personalized feedback generation, manual evaluation indicates that Deepseek outperforms GPT-4o in terms of quality and relevance. We hope that this work encourages greater research attention towards educational applications of AI in low-resource and underserved contexts.
Limitations
While this study provides valuable insights into the use of vision-language and large language models (VLMs and LLMs) for automated assessment in multilingual, low-resource contexts, several limitations should be acknowledged:
Educational Scope
The study was conducted exclusively in Indonesian public elementary schools, specifically in Grade 4 classrooms following the national curriculum (Kurikulum Merdeka). It focused on two subject areas: Mathematics (covering the introductory chapter on fractions) and English (focusing on the topic of parts of the house). As such, the findings may not be generalizable to other subjects, grade levels, or curricula. Geographically, the research was limited to three provinces—West Java (Java Island), West Nusa Tenggara (Lombok Island), and West Sumatra (Sumatra Island)-–which, while diverse, may not fully represent the broader variation in educational contexts across Indonesia or other countries.
Models
The models used in our evaluations include OpenAI’s GPT-4o, Meta’s LLaMA 3.1–70B Instruct, Qwen 2.5–VL–72B Instruct, and DeepSeek Chat. While these models represent the current state of the art, their training data and evaluation strategies are primarily optimized for English and other globally dominant contexts. As a result, they may struggle to fully capture the nuances of student responses written in Bahasa Indonesia.
Acknowledgments
This research was supported by the 2023 SEAMEO-Australia Education Links Award, jointly organized by the Southeast Asian Ministers of Education Organization (SEAMEO) Secretariat and the Australian Government Department of Education.
Ethics Statement
This study strictly adheres to ethical research practices in AI and education:
  • All student answer sheets were anonymized prior to analysis. Identifying information, including names, school names, and class identifiers, was removed to protect student privacy and comply with ethical guidelines for research involving minors.
  • Written informed consent was obtained from school administrators and participating teachers. Participation in the study was voluntary, and students were not penalized for opting out.
  • The inclusion of both urban and rural schools was an intentional decision to ensure representation across socio-economic and educational divides. However, we recognize that the deployment of AI tools in such settings must be approached cautiously to avoid reinforcing existing inequalities. This study advocates for equitable development, localization, and participatory design of AI tools in education, particularly when applied in under-resourced areas.
  • To mitigate risks associated with overreliance on AI outputs, all AI-generated scores and feedback were reviewed by experienced teachers. We emphasize that AI should augment—not replace—human judgment in educational assessment, especially when dealing with young learners.
Appendix A Hyperarameter Setup
We use the following default hyperparameters: temperature = 1.0, top-p = 1.0, and top-k = 1.0 for all tasks, including OCR of student papers, scoring, and generating feedback. The max_tokens parameter is also set to its default to allow the model to generate output without restrictions.
Appendix B Prompts List
Figures 2, 3, and 4 show the prompts we use to generate outputs for the OCR task, score student answers, and provide feedback based on the student’s assignment performance.
Prompt for reading the image (OCR) This is an image of an answer sheet with texts written in either English or Indonesian. Please extract all answers from the image. Adjust the numbering in your response to match the actual number of questions on the answer sheet. Use the following JSON format in your output, and do not output anything else. { ’Nama’: <value>, ’Kelas’: <value>, ’PILIHAN GANDA’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, ’ISIAN’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, ’ESSAY’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, }
Figure 2: Prompt for reading the image (OCR) using LLM
Prompt for scoring The maximum score for this question is {max_score}. Please follow this marking criteria when deciding the score for the student’s answer {marking_criteria} Student answer: {student_answer} Answer key: {gold_answer} What is the appropriate score for the student in a range of 0 and {max_score}? Please only output the score in your response!’
Figure 3: Prompt for scoring using LLM
Prompt for generating the feedback Write in Indonesian a personalised feedback (less than 8 sentences) for a student {student_name} based on the evaluation results over his/her exam answer. Please use this JSON data by focusing on obtained_score and learning_objective. {detailed_feedback}
Figure 4: Prompt for generating the feedback using LLM
Appendix C Human Evaluation Guideline on Personalised Feedback
We evaluate the quality of personalized feedback along four dimensions using a 1–5 rating scale, where 1 indicates the lowest quality and 5 indicates the highest. The four dimensions are Correctness, Personalization, Clarity, and Educational Value / Helpfulness. Correctness assesses whether the feedback is factually accurate based on the student’s response, the answer key, and the rubric. Personalization measures how well the feedback is tailored to the student’s specific answer, including whether it addresses actual strengths, weaknesses, or errors rather than offering generic comments. Clarity evaluates whether the feedback is easy to understand, well-structured, and communicated in an age-appropriate and supportive tone. Educational Value / Helpfulness considers the extent to which the feedback supports learning and encourages the student to reflect and improve. Evaluators are instructed to use these criteria consistently when assigning scores.