### v2

# From Handwriting to Feedback: Evaluating VLMs and LLMs for AI-Powered Assessment in Indonesian Classrooms

Abstract
Despite rapid progress in vision–language and large language models (VLMs and LLMs), their effectiveness for AI-driven educational assessment in real-world, underrepresented classrooms remains largely unexplored. We evaluate state-of-the-art VLMs and LLMs on over 14K handwritten answers from grade-4 classrooms in Indonesia, covering Mathematics and English aligned with the local national curriculum. Unlike prior work on clean digital text, our dataset features naturally curly, diverse handwriting from real classrooms, posing realistic visual and linguistic challenges. Assessment tasks include grading and generating personalized Indonesian feedback guided by rubric-based evaluation. Results show that the VLM struggles with handwriting recognition, causing error propagation in LLM grading, yet LLM feedback remains pedagogically useful despite imperfect visual inputs, revealing limits in personalization and contextual relevance.
1 Introduction
Vision–language models (VLMs) Liu et al. (2023, 2024b); Steiner et al. (2024) and large language models (LLMs) Touvron et al. (2023a); Team (2024); Team et al. (2024); OpenAI et al. (2024) have demonstrated impressive reasoning capabilities Wang et al. (2023); Wei et al. (2022), including solving complex academic tasks such as university-level physics Yeadon and Hardy (2024) and competition-grade mathematics problems Zhang et al. (2024). These advancements have driven growing interest in applying such models to education. Common areas of application include automated grading Chiang et al. (2024), teaching support Hu et al. (2025), feedback generation Morris et al. (2023), and content creation Westerlund and Shcherbakov (2024).
Figure 1: AI-powered assessment using VLM and LLM.
However, most VLM- and LLM-based educational tools have been developed in Global North contexts Lee and Zhai (2025); Yancey et al. (2023), limiting their applicability in classrooms where curricula, languages, and instructional practices differ substantially. Even within these well-resourced settings, only a few studies (e.g., Daniela et al. 2025) have explored end-to-end AI assessment on handwritten student work—and typically on very small samples, such as 24 responses—while the majority of existing benchmarks continue to rely on clean, digital text Chiang et al. (2024).
In Indonesia, paper-based assignments remain the preferred practice in most schools. Many schools, especially in rural areas, lack consistent access to digital devices, highlighting the need for AI systems that function effectively in low-tech environments. Second, using handwritten responses helps reduce the risk of academic dishonesty, such as students who rely on AI tools to generate answers.
Our contributions are as follows:
  • We release the first dataset of handwritten student answer sheets for AI-based educational assessment, comprising over 14K responses collected from six primary schools in Indonesia—three rural and three urban. The dataset covers Grade 4 Mathematics and English, with questions and scoring rubrics designed by experienced teachers. All answers were manually transcribed and graded by professional educators. [^1] To ensure ethical use and protect student privacy, all personally identifiable information (e.g., student names, grade levels, and school names) has been removed. In Indonesia, English is taught as a second language, while Mathematics often involves handwritten explanations in Indonesian (see essay examples in Figure 5).
  • We introduce a multimodal pipeline that integrates vision–language models (VLMs) and large language models (LLMs), as illustrated in Figure 1. We compare several state-of-the-art models for grading student answers and find that GPT-4o with vision input achieves the highest accuracy and feedback quality.
  • We conduct a manual evaluation of LLM-generated feedback in Indonesian and find that, even when based on imperfect input (e.g., OCR errors), the feedback tends to be clear and factually correct. However, personalization and helpfulness remain notable areas of concern.
2 Related Work
Previous studies have investigated the use of LLMs as graders for student assignments and exams. For example, Chiang et al. (2024) used GPT-4 to automatically grade 1,028 student essays in a university-level course titled Introduction to Generative AI. Their findings suggest that LLM-based graders were generally well accepted by students; however, the models occasionally did not follow the grading rubric. In a related study, Yancey et al. (2023) used GPT-3.5 and GPT-4 to score essays in a high-stakes English proficiency test, demonstrating that LLM-generated scores can achieve high agreement with human raters.
Stahl et al. (2024) used Mistral (Jiang et al., 2023) and LLaMA-2 (Touvron et al., 2023b) to assess English student essays and generate feedback, finding that scoring accuracy had limited influence on student’s perceived usefulness of the feedback. Similarly, Morris et al. (2023) applied a Longformer-based language model (Botarleanu et al., 2022) to generate formative feedback on student-written summaries of English textbooks.
Unlike these prior studies, our work focuses on handwritten responses from grade 4 primary school students in Indonesia, covering both English and mathematics. We also evaluate a complete multimodal pipeline that integrates a VLM for handwriting recognition and LLMs for grading and feedback generation—introducing new challenges related to noisy input, multilingual content, and real-world constraints in low-tech, underrepresented classroom settings.
3 Dataset Construction
Assessment Design
We developed assessment instruments for grade 4 primary school students in two subjects: Mathematics and English. The items were designed from scratch based on a thorough analysis of the national curriculum and corresponding learning objectives. Each subject assessment consisted of 10 multiple-choice questions (MCQs), 10 short-answer questions, and 2 essay questions. All items were created by experienced senior subject teachers—an English teacher and a Math teacher—each with over 10 years of classroom experience and a Master’s degree in Education. In addition to writing the assessment items, these teachers developed detailed scoring rubrics for the short-answer and essay questions, as well as answer keys for the MCQs. Standardized answer sheets were also prepared to collect student responses. The assessment and rubric examples for both English and Math subjects are presented in Appendix A.
Data Collection
Data collection was carried out in six primary schools, evenly divided between rural (Sumatra and Nusa Tenggara Islands) and urban (Java Island) settings. Each classroom included approximately 20 to 30 students. Students had up to 30 minutes to complete their answers on a standardized answer sheet.
In total, we collected 646 handwritten answer sheets from these assessments, with a total of more than 14K student answers. Of these, 414 were collected from urban schools and 232 from rural schools (See Figure 5 and Figure 6 in Appendix for example of student answer sheets). The disparity in sample size between urban and rural areas is primarily due to larger class sizes typically found in urban schools compared to their rural counterparts.
English Math
Model M S E Total M S E Total
OCR by GPT4o
GPT4o 2.8 14.6 5.6 11.7 2.3 16.3 1.5 8.2
Llama 3.1 (70B) 2.8 18.7 9.3 14.5 2.3 10.6 27.5 2.2
Qwen2.5 (72B) 2.8 14.9 16.6 14.7 2.3 19.1 5.8 7.1
Deepseek (671B) 2.8 12.6 9.8 11.9 2.3 22.8 6.7 8.1
OCR by Human
GPT4o 0.0 9.2 2.7 7.9 0.0 2.9 5.7 1.5
Llama 3.1 (70B) 0.0 14.4 2.3 11.6 0.0 9.8 19.1 10.3
Qwen2.5 (72B) 0.0 8.4 3.8 9.2 0.0 5.5 8.7 3.3
Deepseek (671B) 0.0 4.4 1.5 6.8 0.0 5.9 8.5 0.8
Table 1: Mean absolute error (MAE) for scoring English and Math, calculated separately for multiple-choice (M), short-answer (S), essay (E), and the total score. Lower values indicate better performance; bolded numbers represent the best results. Scores range from 0 to 100.
4 Experiment
Overall Pipeline
Figure 1 illustrates our pipeline, which begins with a vision–language model (VLM) that performs optical character recognition (OCR) to extract handwritten student responses from scanned answer sheets. The extracted text is then structured into a JSON format and passed to a large language model (LLM), along with the answer key and a teacher-defined rubric. For multiple-choice questions, we apply string matching. For short-answer and essay questions, we run the LLM separately for each question, providing the student’s response, the corresponding answer key, and the assessment rubric. To generate personalized feedback, we provide the LLM with all of the student’s responses, the answer key, the assigned weights, and the rubric.
Model
For OCR, we employ GPT-4o OpenAI et al. (2024), alongside a gold-standard transcription manually prepared by teachers for comaprison. We also evaluated Qwen-VL-72B (see Table 5 in Appendix E) but found its performance suboptimal; hence, GPT-4o was used for all subsequent experiments. For automatic scoring, we compare the performance of GPT-4o, Llama-3.1-Instruct (70B) Touvron et al. (2023b), Qwen2.5-Instruct (72B) Team (2024), and DeepSeek-Chat (671B) Liu et al. (2024a). For personalized feedback generation, we rely on the grading results from GPT-4o, producing two feedback variants using GPT-4o and DeepSeek-Chat, respectively. All prompts and decoding hyperparameters are detailed in Appendix B and Appendix C.
Evaluation
Each answer sheet image was manually transcribed and scored by professional teachers. We compared the LLM-generated scores against these gold-standard scores across three question types: multiple-choice, short-answer, and essay, using mean absolute error (MAE) as the evaluation metric. For personalized feedback, we conducted a manual evaluation covering four aspects—Correctness, Personalization, Clarity, and Educational Value/Helpfulness—rated on a 1–5 scale, where 1 indicates the lowest quality. [^2] This evaluation was carried out by experienced educators with a Master’s degree in teaching. The evaluation guidelines and definitions for each aspect are provided in the Appendix.
Model Correctness Personalization Clarity Helpfulness
English
GPT-4o 4.00 ($\pm$1.0) 3.96 ($\pm$0.7) 3.64 ($\pm$0.8) 3.60 ($\pm$0.8)
Deepseek 3.96 ($\pm$0.9) 3.88 ($\pm$0.8) 4.04 ($\pm$0.9) 3.96 ($\pm$1.1)
Math
GPT-4o 3.84 ($\pm$0.9) 3.72 ($\pm$0.9) 3.92 ($\pm$1.2) 3.68 ($\pm$0.8)
Deepseek 3.88 ($\pm$0.3) 2.96 ($\pm$0.5) 4.00 ($\pm$0.0) 2.92 ($\pm$0.3)
Table 2: Human evaluation by expert teachers on personalized feedback, using a rating scale from 1 to 5, where 1 indicates the lowest score.
5 Result and Analysis
Main Result
Table 1 presents the performance of the LLMs selected in three types of questions: multiple choice, short answer, and essay. When using GPT-4o to extract student responses via OCR, we observe that most model-generated scores are generally competitive. Among them, GPT-4o produces scores that align most closely with human grading for essay questions, achieving the lowest MAE in both English (5.6) and Math (1.5). In contrast, LLaMA-3.1–70B and Qwen-2.5–72B are less reliable, with scores deviating more significantly from human judgments. Short-answer questions remain the most challenging to evaluate: even the best-performing model in this category, LLaMA-3.1-7B for Math, still shows a relatively high MAE of 10.6, indicating a notable gap from human-level accuracy.
However, the results differ when human effort is involved in the OCR task. Most scores become better overall, with Deepseek-chat and GPT-4o emerging as the top-performing models. Deepseek-chat shows strong performance in English (MAE of 4.4 for short answers and 1.5 for essays), while GPT-4o performs best in Math, with only a 2.9 difference in short answers and 5.7 in essays. It is worth noting that MCQ scores remain at 0, as basic string matching is sufficient due to the exact nature of the answers. The impact of OCR performance on LLM scoring is further discussed in Section 5 of OCR performance analysis.
Human Evaluation on Personalised Feedback
Table 2 presents the results of a human evaluation on personalized feedback quality, rated by expert teachers across four dimensions: Correctness, Personalization, Clarity, and Helpfulness (scale 1–5, with scores below 3 considered poor). For English, GPT-4o slightly outperforms Deepseek in correctness and personalization, while Deepseek leads in clarity and helpfulness. In Math, Deepseek shows strong clarity and correctness but performs poorly in personalization and helpfulness, with both scores falling below 3. GPT-4o, on the other hand, maintains more balanced performance across all dimensions.
English Math
Model M S E Total M S E Total
Urban
GPT4o 0.0 2.4 7.2 0.8 0.0 5.8 7.6 2.4
Llama 3.1 (70B) 0.0 7.7 2.9 2.7 0.0 10.3 30.0 10.4
Qwen2.5 (72B) 0.0 1.9 1.3 0.5 0.0 7.6 10.7 3.9
Deepseek (671B) 0.0 1.3 3.5 1.5 0.0 5.6 9.9 1.0
Rural
GPT4o 0.0 21.2 5.2 23.1 0.0 2.5 2.2 0.3
Llama 3.1 (70B) 0.0 26.1 11.4 26.9 0.0 8.8 23.1 9.7
Qwen2.5 (72B) 0.0 19.8 12.5 24.3 0.0 1.7 5.0 2.1
Deepseek (671B) 0.0 14.2 10.1 21.2 0.0 6.4 5.9 0.6
Table 3: MAE for scoring English and Math across urban and rural settings, calculated separately for multiple-choice (M), short-answer (S), essay (E), and total scores. The OCR results used in this analysis were obtained through human transcription. Lower values indicate better performance; bolded values represent the best results. Each component is scored on a 0–100 scale.
Urban vs. Rural Performance Analysis
Given the significant educational disparities between rural and urban areas, we evaluated the performance of the model in these two settings. To isolate the analysis of LLM scoring capabilities, we use only the human-transcribed OCR results, eliminating recognition errors.
Table 3 presents the MAE scores for English and Math, separated by question type: multiple choice (M), short answer (S), essay (E), and total scores. The results indicate that English MAEs are generally higher in rural settings than in urban settings across all models. For example, GPT-4o achieves a total MAE of only 0.8 in urban English, but this rises sharply to 23.1 in the rural setting. This discrepancy suggests that LLMs may struggle more in interpreting free-form responses from rural students, possibly due to variations in writing style and grammar. In contrast, MAEs for Math tend to be slightly lower in rural areas, although the differences are less pronounced. This may be attributed to the nature of Math questions, which often involve numerical reasoning and have more deterministic answers, reducing ambiguity in scoring.
OCR Performance Analysis
Given the differences in MAE between the GPT-4o OCR outputs and human transcription shown in Table 1, we further analyze the OCR performance of GPT-4o and evaluate the extent to which recognition errors propagate to the subsequent scoring. For this analysis, we use exact string matching to assess accuracy on multiple-choice and short answer questions, and compute ROUGE-L Lin (2004) scores to compare GPT-4o and human transcriptions for essay questions.
Table 4 shows that the OCR performance is generally higher for English than for Math. Within English, responses from urban students yield higher exact match and ROUGE-L scores compared to those from rural students, possibly due to differences in handwriting clarity or writing conventions. For Math, the OCR accuracy is overall lower than that of English, but the performance gap between urban and rural settings is less pronounced. This suggests that while English responses may be more affected by region-specific handwriting variability, Math responses, often more structured and numerical, are comparatively stable across regions.
English Math
Area EM(M) EM(S) RL(E) EM(M) EM(S) RL(E)
Urban 82.1 67.1 60.3 62.3 23.3 21.0
Rural 71.7 61.8 60.1 62.5 27.9 24.8
All 78.5 65.3 60.2 62.4 24.9 22.3
Table 4: OCR-based performance (GPT-4o) across Urban, Rural, and All settings for English and Math: EM = exact match, RL = ROUGE-L F1, M = multiple choice, S = Short Essay, E = Essay.
6 Conclusion
We release the first dataset of handwritten student answer sheets for AI-based assessment in Indonesia, covering over 14K Math and English responses from rural and urban schools. Using this dataset, we evaluate VLMs and LLMs for grading and personalized feedback, finding that GPT-4o and DeepSeek (671B) perform competitively with teacher scores, with DeepSeek producing more contextually relevant feedback. This work lays the foundation for AI-powered assessment in low-resource and underrepresented educational contexts.
Limitations
While this study provides valuable insights into the use of vision-language and large language models (VLMs and LLMs) for automated assessment in multilingual, low-resource contexts, several limitations should be acknowledged:
Educational Scope
The study was conducted exclusively in Indonesian public elementary schools, specifically in Grade 4 classrooms following the national curriculum (Kurikulum Merdeka). It focused on two subject areas: Mathematics (covering the introductory chapter on fractions) and English (focusing on the topic of parts of the house). As such, the findings may not be generalizable to other subjects, grade levels, or curricula. Geographically, the research was limited to three provinces—West Java (Java Island), West Nusa Tenggara (Lombok Island), and West Sumatra (Sumatra Island)-–which, while diverse, may not fully represent the broader variation in educational contexts across Indonesia or other countries.
Models
The models used in our evaluations include OpenAI’s GPT-4o, Meta’s LLaMA 3.1–70B Instruct, Qwen 2.5–VL–72B Instruct, and DeepSeek Chat. While these models represent the current state of the art, their training data and evaluation strategies are primarily optimized for English and other globally dominant contexts. As a result, they may struggle to fully capture the nuances of student responses written in Bahasa Indonesia.
Acknowledgments
Reserved due to double-blinds.
Ethics Statement
This study strictly adheres to ethical research practices in AI and education:
  • All student answer sheets were anonymized prior to analysis. Identifying information, including names, school names, and class identifiers, was removed to protect student privacy and comply with ethical guidelines for research involving minors.
  • Written informed consent was obtained from school administrators and participating teachers. Participation in the study was voluntary, and students were not penalized for opting out.
  • The inclusion of both urban and rural schools was an intentional decision to ensure representation across socio-economic and educational divides. However, we recognize that the deployment of AI tools in such settings must be approached cautiously to avoid reinforcing existing inequalities. This study advocates for equitable development, localization, and participatory design of AI tools in education, particularly when applied in under-resourced areas.
  • To mitigate risks associated with overreliance on AI outputs, all AI-generated scores and feedback were reviewed by experienced teachers. We emphasize that AI should augment—not replace—human judgment in educational assessment, especially when dealing with young learners.
Appendix A Example of Assessment Items and Rubrics
Example assessment items for English and Math are shown below. The questions were designed in accordance with the Indonesian curriculum to ensure that students can understand the questions given.
  • Math
    (Multiple choice) Pecahan mana yang paling besar? (Which fraction is the largest?) A. 1/9 B. 1/8 C. 1/7 D. 1/6
    (Essay) Terdapat empat buah pecahan: 3/10, 8/10, 1/10, 6/10. Urutkan pecahan-pecahan tersebut dari yang paling kecil ke yang paling besar (There are four fractions: 3/10, 8/10, 1/10, 6/10. Sort the fractions from smallest to largest)
  • English
    (Multiple choice) Look at the picture and choose the correct word for it [Picture is provided] A. bathroom; B. living room; C. Dining room
    (Essay) Look at the picture. [Picture is provided] Write a text, consisting of information: How many rooms are there? (write 1 sentence) What are the rooms? (write 1 sentence) How is the position of the rooms? (Write at least 5 sentences)
Moreover, we also provide an example of a rubric used to score the essay here. The rubric was developed by a teacher with over 10 years of classroom experience, ensuring it is accurate and aligned with the assessment criteria.
  • Math (written in Indonesian)
    Score 0: Jawaban ada dan benar, tetapi tanpa alasan (The answer is correct, but without a reason)
    Score 1: Jawaban ada dan benar, hanya menyebutkan 1/10 lebih besar dari 1/100. (The answer is correct, it just states that 1/10 is greater than 1/100.)
    Score 2: Jawaban ada dan benar, hanya menyebutkan 10 lebih kecil dari 100. (The answer is correct, it only states that 10 is less than 100.)
    Score 3: Jawaban ada dan benar, menyebutkan karena pecahan tersebut adalah pecahan dengan pembilang satu sehingga semakin kecil penyebut, semakin besar pecahannya. (The answer is correct, stating that the fraction is a fraction with a numerator of one, so the smaller the denominator, the larger the fraction.)
  • English (written in English)
    Score 0: The response does not meet the above criteria (e.g., incorrect sentence structure, incorrect use of prepositions, or no response).
    Score 1: The response is an incomplete sentence, but uses the correct preposition of place. Example: "bedroom under the living room."
    Score 2: The student’s response is a complete sentence but has 1 or 2 punctuation errors. Example: "The bedroom is under living room."
    Score 3: The student’s response is a complete sentence with correct grammar and punctuation. Example: "The bedroom is under the living room."
Appendix B Hyperarameter Setup
We use the following default hyperparameters: temperature = 1.0, top-p = 1.0, and top-k = 1.0 for all tasks, including OCR of student papers, scoring, and generating feedback. The max_tokens parameter is also set to its default to allow the model to generate output without restrictions.
Appendix C Prompts List
Figures 2, 3, and 4 show the prompts we use to generate outputs for the OCR task, score student answers, and provide feedback based on the student’s assignment performance.
Prompt for reading the image (OCR) This is an image of an answer sheet with texts written in either English or Indonesian. Please extract all answers from the image. Adjust the numbering in your response to match the actual number of questions on the answer sheet. Use the following JSON format in your output, and do not output anything else. { ’Nama’: <value>, ’Kelas’: <value>, ’PILIHAN GANDA’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, ’ISIAN’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, ’ESSAY’: { ’1’: <value>, ’2’: <value>, // Adjust numbering based on the answer sheet}, }
Figure 2: Prompt for reading the image (OCR) using LLM
Prompt for scoring The maximum score for this question is {max_score}. Please follow this marking criteria when deciding the score for the student’s answer {marking_criteria} Student answer: {student_answer} Answer key: {gold_answer} What is the appropriate score for the student in a range of 0 and {max_score}? Please only output the score in your response!’
Figure 3: Prompt for scoring using LLM
Prompt for generating the feedback Write in Indonesian a personalised feedback (less than 8 sentences) for a student {student_name} based on the evaluation results over his/her exam answer. Please use this JSON data by focusing on obtained_score and learning_objective. {detailed_feedback}
Figure 4: Prompt for generating the feedback using LLM
Appendix D Human Evaluation Guideline on Personalised Feedback
We evaluate the quality of personalized feedback along four dimensions using a 1–5 rating scale, where 1 indicates the lowest quality and 5 indicates the highest. The four dimensions are Correctness, Personalization, Clarity, and Educational Value / Helpfulness. Correctness assesses whether the feedback is factually accurate based on the student’s response, the answer key, and the rubric. Personalization measures how well the feedback is tailored to the student’s specific answer, including whether it addresses actual strengths, weaknesses, or errors rather than offering generic comments. Clarity evaluates whether the feedback is easy to understand, well-structured, and communicated in an age-appropriate and supportive tone. Educational Value / Helpfulness considers the extent to which the feedback supports learning and encourages the student to reflect and improve. Evaluators are instructed to use these criteria consistently when assigning scores.
Appendix E Qwen-VL-72B OCR Results and The Decision of Using GPT-4o
For OCR, we decided to use GPT-4o instead of another open-source LLM, as we had already conducted a comparison between GPT-4o and Qwen-VL-72B across two subjects: English and Math. The results for Qwen-VL-72B are shown in Table 5, with an average English score of 67.5% and a Math score of 35.1%, which are both slightly lower than GPT-4o’s performance (English: 68%, Math: 36.5%). Therefore, we chose GPT-4o for the OCR tasks.
English Math
Area EM(M) EM(S) RL(E) EM(M) EM(S) RL(E)
Urban 80.9 67.2 59.3 59.9 23.3 20.5
Rural 73.8 60.2 60.2 60.1 21.8 23.8
All 78.5 64.7 59.3 61.0 22.8 21.6
Table 5: OCR-based performance (Qwen-VL-72B) across Urban, Rural, and All settings for English and Math: EM = exact match, RL = ROUGE-L F1, M = multiple choice, S = Short Essay, E = Essay.
Appendix F Additional Data Details: Students’ Average Scores
For completeness, we also present the dataset information, including the average scores for each subject and question type. The students’ average scores, manually graded by humans, are shown in Table 6. Based on these scores, we calculate the mean absolute error (MAE) for each model, as presented in Table 1, with English and Math evaluated separately according to each model’s grading performance.
English Math
Area M S E M S E
Urban 85.27 63.29 30.90 52.00 67.43 61.36
Rural 72.29 41.21 26.60 34.80 63.38 66.71
All 80.53 55.22 30.90 52.00 67.43 61.36
Table 6: The students’ average scores, manually graded by humans. M = multiple choice, S = Short Essay, E = Essay.
Appendix G Example Qualitative Analysis
We also present examples of students’ assessments, specifically from rural areas, as shown in Figures 5 and 6. Based on the OCR results, certain patterns indicate that the OCR model (e.g., GPT-4o) struggles to interpret some student handwriting, even though it performs well on multiple-choice questions.
In the case of the math subject, the OCR model occasionally misinterprets numerals. For example, reading a 9 in the denominator as an 8, or incorrectly extracting 1/3 as 1/5. For the English subject, most errors occur in the essay section. The OCR model sometimes misreads words, such as interpreting “bathroom” as “bedroom.” Additionally, the model often fails to predict the answer to the final English question correctly because some students write more responses than the paper allows (a common occurrence in real-life classroom settings), which confuses the OCR system.
Figure 5: Example of a student’s math assessment answer sheet from a rural area. The student’s identity has been obscured for privacy.
Figure 6: Example of a student’s English assessment answer sheet from a rural area. The student’s identity has been obscured for privacy.